Find the curvature, the radius and the center of curvature at a point.



x = t ^ 2 , y = t ^ 3 ; t = 1/2


1
Expert's answer
2021-12-07T10:03:07-0500
"K=\dfrac{|x'y''-x''y'|}{[(x')^2+(y')^2]^{3/2}}""x'_t=6t, x''_{tt}=6""y'_t=3t^2-3,y''_{tt}=6t-3""K=\dfrac{|6t(6t-3)-6(3t^2-3)|}{[(6t)^2+(3t^2-3)^2]^{3/2}}"

"t=2"

"K=\dfrac{|6(2)(6(2)-3)-6(3(2)^2-3)|}{[(6(2))^2+(3(2)^2-3)^2]^{3/2}}=\dfrac{2}{125}""=0.016"

The radius of curvature of a curve at a point is called the inverse of the curvature "K" of the curve at this point:



"R=\dfrac{1}{K}=62.5""y'_x=\dfrac{y'_t}{x'_t}=\dfrac{3t^2-3}{6t}=\dfrac{t^2-1}{2t}""(y'_x)'_t=\dfrac{2t^2-t^2+1}{2t^2}=\dfrac{t^2+1}{2t^2}""y''_{xx}=\dfrac{t^2+1}{2t^2(6t)}=\dfrac{t^2+1}{12t^3}""x_C=x-\dfrac{y'(1+(y')^2)}{y''}""y_C=y+\dfrac{1+(y')^2}{y''}"

"t=2"



"x=3(2)^2=12, y=(2)^3-3(2)=2""y'=\dfrac{(2)^2-1}{2(2)}=\dfrac{1}{4}""y''=\dfrac{(2)^2+1}{12(2)^3}=\dfrac{5}{96}""x_C=12-\dfrac{\dfrac{1}{4}(1+(\dfrac{1}{4})^2)}{\dfrac{5}{96}}=6.9""y_C=2+\dfrac{1+(\dfrac{1}{4})^2}{\dfrac{5}{96}}=32.4"

"C(6.9, 32.4)"